OpenStudy (anonymous):

if there is a 60% chance of rain tomorrow and a 40% chance of wind and rain, what is the probability that it is windy, given that it is rainy. Round your answer to the nearest percent. A:33% B67% C:71% D:78%

OpenStudy (love.kat123):

i shall help u XD

OpenStudy (anonymous):

thank you

OpenStudy (love.kat123):

hold on a sec

OpenStudy (love.kat123):

i would say A not sure tho

OpenStudy (anonymous):

\[P(A | B) =\frac{ P(A \cap B) }{ P(B) }\]

OpenStudy (anonymous):

pgpilot do you agree with A?

OpenStudy (anonymous):

if B is the event that it is rainy and A is the event that it is windy then \[P(\text{Windy|Rainy})=\frac{ P(\text{Windy AND Rainy}) }{ P(\text{Rainy}) }\]

OpenStudy (anonymous):

then you substitute the values?

OpenStudy (anonymous):

no... .4/.6 = 2/3 = .66666666....

OpenStudy (anonymous):

so its 67%

OpenStudy (anonymous):

yep. conditional probability. time to learn it!

OpenStudy (anonymous):

im on the lesson right now. stupid apex doesn't explain things very well. thanks

OpenStudy (love.kat123):

well its a good thing u didnt pick A sorry i was wrong

OpenStudy (anonymous):

here's a link that may help... https://onlinecourses.science.psu.edu/stat414/node/10

OpenStudy (anonymous):

it does a pretty good job.

OpenStudy (anonymous):

there also other probability topics you can review as well

OpenStudy (anonymous):

ty bb